Amythehedgehog:

Look at this prism and its net. Rectangular prism with length labeled 5 meters, width labeled 7 meters, and height labeled 12 meters. Net of the same rectangular prism with the same dimensionsn labeled. What is the surface area of this rectangular prism?

Amythehedgehog:

1 attachment
Hero:

@Amythehedgehog do you remember the formula for area of a rectangle?

callielovesyhuu:

A=2(wl+hl+hw) plug it in

Hero:

As I was trying to explain earlier we are given the length, width, and height of a rectangular prism and asked to find the surface area of the rectangular prism. Understand that the surface area of a prism is the total area of all the sides of the prism. The rectangular prism is made up of six rectangular sides. There is an area type associated with each pair of opposite sides of the prism: \(A_1 = lw\) \(A_2 = wh\) \(A_3 = lh\) To compute the total area of the sides of the prism, we would sum the areas of each rectangular side: (Note: opposite sides of the prism have the same value) \(A_t = lw + lw + wh + wh + lh + lh\) \(A_t = 2lw + 2wh + 2lh\) \(A_t = 2(lw + wh + lh)\) Since the total area of the sides of the prism gives the surface area of the prism we end up with the following formula: \(SA_\text{ prism} = 2(lw + wh + lh)\) As given in the problem: l = 5 w = 7 h = 12 So you would simply insert the given values to the formula for the Surface area of a prism, then simplify the right hand side. If you need any help doing this @Amythehedgehog let me know.

Amythehedgehog:

I AM CONFUSED

Hero:

Understanding comes by taking the time to read carefully and digest each step mentally. When you do not understand something, the proper reaction is to ask a question about the particular concept or step in the problem solving process you are having trouble processing mentally. Saying "I am confused" isn't exactly the best way to communicate your lack of understanding @Amythehedgehog.

Amythehedgehog:

I am still confused about how to do it, I just got the answer from brainly.com

Amythehedgehog:

I like that website

Hero:

I can help you understand how to do it, but you have to ask questions about what you don't understand.

Hero:

You say you like brainly, however, they gave you the answer yet you still don't understand how to do it.

Amythehedgehog:



Hero:

What is that supposed to be?

Hero:

Math Killed You?

Amythehedgehog:

yes I suck at it and now I have brain farts

Amythehedgehog:

thx for trying tho

Hero:

I think the problem is, you don't trust me to help you understand it.

Amythehedgehog:

I just dont like really long responses thats all, and I have 30 min left to finish 5 tests

Hero:

I only posted that because you weren't here.

Amythehedgehog:

my parents are strict when it comes to grades

Hero:

Had you been here, I would have went over it with you step by step.

Amythehedgehog:

I know u did thats why I went to Brainly.com for ANSWERS.

Hero:

You seem to have zero patience.

Amythehedgehog:

sorry I really dont have time to be patient